OpenStudy (anonymous):

Concentrated hydrochloric acid has a density of 1.15g cm^-3 and contains 30.0% by mass hydrogen chloride. How do you determine the concentration of the hydrochloric acid? What volume of this must be diluted to 5.00 dm3 to give a solution of concentration 0.200 mol dm-3

OpenStudy (aaronq):

you can work backwards. find how many moles there are in 5 L of 0.2 M HCl, use \(C_{Molarity}=\dfrac{n_{solute}}{L_{solution}}\) then how the mass of those moles, use \(n=\dfrac{m}{Molar\;mass}\) since only 30% is HCl by mass, find the actually weight you need, use a ratio. when you find the total mass, find the volume you need, \(\rho=\dfrac{m}{V}\) ..or you can first determine the concentration (molarity) of the solution (find the mass of 1 \(L_{solution}\), 30% of that is \(HCl_{(s)}\) convert it to moles and that's the molarity). Then, use \(C_1V_1=C_2V_2\).